OpenStudy (pokemon23):

The product of two consecutive integers is three less than three times their sum.Find the integers.

OpenStudy (anonymous):

are the numbers 6 and 5

OpenStudy (dumbcow):

n(n+1) = 3(2n+1) -3

Directrix (directrix):

Let x and x + 1 be the consecutive integers x(x+1) = 3(x + x + 1) – 3 x^2 + x = 3(2x + 1) – 3 x = 0 and x = 5 ?

OpenStudy (dumbcow):

n^2 -5n = 0

OpenStudy (pokemon23):

then we factor right

OpenStudy (dumbcow):

right pull out an n

OpenStudy (pokemon23):

n(n-5)=0

OpenStudy (pokemon23):

n=0

OpenStudy (pokemon23):

n=5

OpenStudy (pokemon23):

yup ;D

Directrix (directrix):

So, the pairs are 0,1 and 5,6

OpenStudy (pokemon23):

how is it 0 and 1?

OpenStudy (pokemon23):

I understand how it 0 where did the 1 come from

OpenStudy (pokemon23):

Nvm now I know why its 0 and 1 because it consecutive integers going by one silly me